# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

### v2

@@ 4. PinRec
In this section, we discuss the architecture, methodology, and serving strategies for the proposed recommender system, PinRec. Leveraging multi-token, autoregressive generation with outcome conditioning, we present an expressive, diverse, and performant transformer-based generative retrieval technique. We further discuss how PinRec is served in real-time through an in-house implementation of GPT-2 powered by NVIDIA Triton.
4.1. Representation Learning for Items
A critical prerequisite for effective optimization of a sequential recommendation system is an effective input and output representation for the heterogeneous set of items. We denote the representation of some item $i\in\mathcal{I}$ of type $T_{i}=T(i)$ as $f_{T_{i}}(i)$, where there exist representation models $f_{t}:\mathcal{F}(t)\to\mathcal{R}$ for each item type $t\in\mathcal{T}$ mapping the input feature space to the chosen representation space $\mathcal{R}$. For our purposes, we consider two item types ($\mathcal{T}$): Pins and search queries, which represent two core entities at Pinterest.
While some work has often leveraged ID-based representation spaces (Singh et al., 2024), recent work by Liu et al. (2024) and Yang et al. (2024) demonstrates that semantic IDs can lead to poor performance due to representational collapse. To compound onto this, given a larger set of item types $|\mathcal{T}|$, it may be challenging to optimize representations across many different types in a compatible representational space. On the contrary, Agarwal et al. (2024) observe that real-valued vector representations (i.e., $\mathcal{R}\subseteq\mathbb{R}^{d}$) generalize across item types.
[…]
As input to each embedder, we featurize the item $i$ given the chosen set of features $\mathcal{F}(T(i))$ and concatenate the features into a vector $\mathbf{v}_{i}$. To featurize search queries, we use the trained OmniSearchSage query representation (Agarwal et al., 2024), and to featurize Pins, we use the trained PinSage representation (Ying et al., 2018). Given that Pins are the core entity of Pinterest, we choose to implement a large, sharded set of hash-based embedding tables to store a learned ID embedding for each item. The output of the input embedder for item $i$ is a real-valued, $L_{2}$ normalized embedding $\mathbf{i}=f_{t}(\mathbf{v}_{i})$.
4.2. Sequential Modeling Architecture
To construct a sequential recommendation system, we leverage a causal transformer decoder architecture. Provided a sequence of input embeddings, we repeatedly apply multi-head attention followed by feedforward layers. Critically, we apply a causal attention mask at each token, i.e., preventing tokens at time $t$ from attending to any token at $t^{\prime}>t$. The output of the transformer $f_{\theta}(H(u,t_{\mathrm{max}}))$ is a sequence of hidden states from the last layer, and we apply an output head to obtain representations of recommended items.
We base our transformer decoder architecture on the GPT-2 model. We use an in-house implementation that optimizes existing solutions (e.g., HuggingFace [^1] https://huggingface.co/openai-community/gpt2), while allowing for additional functionality. For example, we incorporate the surface for each user interaction and the raw timestamp, following Pancha et al. (2022).
To generate the output representations for item recommendation, we construct an output head $O$ as a small MLP. While the output head is typically conditioned on the transformer’s hidden state, its input can incorporate other forms of conditioning. To that end, we denote any “auxiliary conditioning” as $\Phi(H(u,t_{\mathrm{max}}))$, and we explore this further in Section 4.3.1. We normalize the outputs so they are in the same representational space as the item representations $\mathbf{i}$. To summarize, the output representation $\hat{\mathbf{i}}_{u,t}$ at timestep $t$ for user $u$, applies the output head to the transformer’s last hidden state and any auxiliary conditioning (Equation 1).
(1) $\hat{\mathbf{i}}_{u,t}=O(f_{\theta}(H(u,t_{\mathrm{max}})),\Phi(H(u,t_{\mathrm{max}})))$
4.3. Optimizing Item Recommendations
In this section, we detail our approach to optimizing our item representations with a sequential objective. The standard choice to learn a distribution over personalized item recommendations is simply to employ a “next-token” prediction objective. Beyond directly taking into account the ground-truth user activity and engagement distribution, the next-token objective is effective, efficient, and ubiquitous. Following Pancha et al. (2022), we employ a generic sampled softmax loss to optimize our output item representations against our next item targets, using in-batch and random negatives, and with temperature scaling. Since in-batch negatives are sampled from a biased distribution based on our sequences, we employ sample probability correction through estimating the conditional frequency of each of the items $Q(\mathbf{i}_{c})=P(\mathbf{i}_{c}\mid\hat{\mathbf{i}}_{u,t})$ (Bengio and Senecal, 2008) with a count-min sketch (Cormode and Muthukrishnan, 2005).
Based on the vector representations of the output item representation $\hat{\mathbf{i}}_{u,t}$ and a candidate item $\mathbf{i}_{c}$, we define their similarity function $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{\mathbf{i}_{c}})$, applying a correction to account for sampling bias. The sampled softmax loss defined in Equation 2 maximizes the similarity of our output item representation $\hat{\mathbf{i}}_{u,t}$ and the next item representation $\mathbf{i}_{u,t+1}$, relative to in-batch and random negative examples $\mathbf{N}$.
(2) $L_{\mathrm{s}}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1})=-\log\left(\frac{\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1}))}{\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1}))+\sum_{\mathbf{i}_{n}\in\mathbf{N}}\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{n}))}\right)$
Our empirical next item loss function, over a user sequence $H(u,t_{\mathrm{max}})$ and sampled negatives $\mathbf{N}$, maximizes next item similarity across the entire sequence, as shown in Equation 3 (Pancha et al., 2022).
(3) $L_{\theta}^{\mathrm{next}}(H(u,t_{\mathrm{max}}))=\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t}L_{s}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1})$
To optimize the model parameters $\theta$ and consequently learn an optimal set of item recommendations, we minimize the empirical loss across user sequences and sampled negatives in our dataset $\mathcal{D}$.
(4) $\arg\min_{\theta}\mathbb{E}_{H(u,t_{\mathrm{max}})\sim\mathcal{D}}[L_{\theta}^{\mathrm{next}}(H(u,t_{\mathrm{max}}))]$
4.3.1. Outcome-Conditioned Generation
Although we optimize with respect to item targets conditioned on the user’s history, the next item loss provides no additional contextual information for the output representation by itself. For instance, given a user who has a propensity for low-effort engagement, the standard next token objective tends to generate recommendations that reinforce this same behavior. Though prior approaches leverage outcome-based loss weighting to resolve this, it requires domain knowledge and nevertheless requires significant amounts of tuning to achieve the desired outcome for each user. Alternatively, RL approaches are fraught with training instability and/or inefficiency and a significant need for hyperparameter tuning (Schulman et al., 2017). On the contrary, we propose reframing this problem by simply conditioning the generation of item recommendations directly on the desired outcomes (i.e., outcome or goal-conditioning).
To that end, we additionally condition the output head $O$ of the model on the desired outcome. To enable representing different outcomes, we employ a set of learnable embeddings for each possible action outcome (e.g., user repins, user clicks, etc.). By conditioning the output generation on a representation of the desired outcome (i.e., $\Phi(H(u,t_{\mathrm{max}}))$ represents $a_{\mathrm{desired}}\in\mathcal{A}$), we can generate outcome-specific, personalized item representations $\hat{\mathbf{i}}_{u,t}$.
At training time, we leverage user sequences to optimize the output representation with respect to the targets. Since we cannot dictate the desired outcome in offline data post-hoc, we optimize with respect to outcomes achieved in the data (Ghosh et al., 2021). Hence, we can define the auxiliary conditioning during training time as $\Phi(H(u,t_{\mathrm{max}}))=a(i_{t+1})$, corresponding to the target item $i_{t+1}$.
During inference, we can specify the desirability of each outcome and condition our generation according to these preferences. When the system is deployed, this allows us to induce these desirable outcomes from users. Formally, we define a per-action budget distribution $\mathcal{B}:\mathcal{A}\to\mathbb{R}$ that maps each action to a non-negative budget such that $\sum_{a\in\mathcal{A}}\mathcal{B}(a)=1$. Hence, as the budget for any action grows, it becomes more desirable. Based on the budget $\mathcal{B}$ and the desired numbers of recommended items $N$, we can either deterministically or stochastically allocate a fixed number of items for each action during generation (e.g., based on the expectation, each action $a$ can deterministically be allocated $\mathcal{B}(a)\cdot N$ items).
4.3.2. Windowed Multi-Token Generation
One of the principal assumptions in the next token objective is that there is strict ordering between the elements of the sequence. However, that reasoning is not necessarily applicable to social media platforms, where a user may engage with suggestions on their feed in an arbitrary order within a session. Moreover, a user may engage with certain items later on within a session (or perhaps at the very end) that they may have come across at the very beginning of the session. The next token objective simply ignores these crucial facets of recommendation systems for platforms such as Pinterest.
Motivated by the absence of strict ordering in sequential user engagement, we extend the next-token prediction objective. We assume that there exists a timestep window of size $\Delta$ within which user activity is roughly time-equivalent (i.e., their order within that window is not significant). To address the latter issue of delayed user engagement of an item relative to the current timestep, we employ an efficient multi-token objective which enables prediction beyond the next token at subsequent timesteps $t^{\prime}\geq t+1$.
Often, multi-token prediction for language models involves predicting a handful (e.g., 5-10) of successive tokens at timesteps $t+1$ to $t+k$ after the current token at timestep $t$ (Gloeckle et al., 2024). However, this leads to a similar issue of penalizing any predictions that do not match the target at exactly some timestep $t^{\prime}\in[t+1,t+k]$, even if a matching target may exist within the window $[t+1,t+\Delta]$. To incorporate the time-equivalent window size $\Delta$, we propose a windowed multi-token objective. That is, for any future timestep window of size $\Delta$ starting at some timestep $t^{\prime}>t$, all items that are engaged within that window are targets for our prediction task.
We choose to condition the output generation at timestep $t$ on both the target timestep $t^{\prime}>t$ and its action outcome (though it can work without outcome-conditioning). That is, we extend $\Phi(H(u,t_{\mathrm{max}}))$ to represent $\{t^{\prime},a(i_{t^{\prime}})\}$ at training time. Similarly to outcome-conditioning, we can decide $t^{\prime}$ (and the corresponding desired outcomes based on $\mathcal{B}$) at inference time as desired.
Our mathematical formulation for the windowed multi-token objective is shown in Equation 5, where $p_{t^{\prime}}$ denotes the chosen distribution over target timestep $t^{\prime}$ given $t$ (e.g., a uniform distribution from $[t+1,t_{\mathrm{lim}}]$ for some $t_{\mathrm{lim}}$). Note that a non-zero likelihood of sampling $t^{\prime}=t+1$ additionally enables this multi-token objective to perform next-token prediction and autoregressive generation.
(5) $\begin{split}L_{\theta}^{\mathrm{mt}}&(H(u,t_{\mathrm{max}}))=\\ &\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t}\mathbb{E}_{t^{\prime}\sim p_{t^{\prime}}(\cdot\mid t)}\Big{[}\min\{L_{s}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,i})\mid\forall i\in[t^{\prime},t^{\prime}+\Delta]\}\Big{]}\end{split}$
Intuitively, the optimization problem associated with the windowed multi-token objective can be posed as a question: can we sequentially predict any future items from the set of items that the user will engage with between timestep $t^{\prime}$ and $t^{\prime}+\Delta$? We find that this optimization is more intuitively aligned with the properties of user interaction on Pinterest compared to next-token prediction, while allowing for significantly reduced serving latency due to multiple items being predicted at a single token.

### v4

@@ 4. PinRec
In this section, we describe the model architecture and serving infrastructure for PinRec. We start with an overview of the model, including outcome-conditioned generation and multi-token prediction. Next, we explain how items in the sequence are encoded and processed, and how the training loss is computed. Finally, we describe inference mode operation and provide details about our real-time serving architecture, which is powered by Nvidia Triton.
4.1. Model Architecture
We use a transformer decoder architecture as the backbone of our PinRec model. The user interaction history is first featurized and fed into a stack of transformer layers along with learned position embeddings to capture temporal ordering. At each layer, we apply causal attention masking to ensure that tokens at time $t$ can only attend to tokens at positions $t^{\prime}\leq t$, preventing information leakage from future interactions. The stack of $L$ transformer layers outputs a sequence of contextualized hidden states: $\mathbf{h}_{u,t}=\text{TransformerStack}^{(L)}(\mathbf{x}_{u,1:t})$.
where $\mathbf{x}_{u,1:t}$ represents the featurized user interaction sequence up to timestep $t$, and $\mathbf{h}_{u,t}$ is the resulting hidden state at position $t$ from the final layer. These hidden states are then processed through an output head $O$ to generate the next item predictions.
Outcome-Conditioned Generation. Unlike language modeling, where the output sequence receives a single reward value (PPO22), users interact with recommended items through diverse engagement types—saving, sharing, commenting, or clicking. To effectively optimize across these varied engagement metrics, generative retrieval methods must account for this behavioral diversity (MultiBehaviorSASRec24). Furthermore, practical deployment often requires specifying the target surface or context for the next recommendation, even when the user’s historical sequence spans multiple surfaces.
To address these requirements, we design our model to accept conditioning signals at each generation step, enabling dynamic control over model behavior during inference. Specifically, we condition the output head $O$ on a desired set of outcomes, allowing the model to generate targeted recommendations. We implement this through learnable embeddings for each possible outcome type (e.g., engagement actions, surface constraints). By conditioning the generation process on these outcome representations, our model produces outcome-specific, personalized item representations:
(1) $\hat{\mathbf{i}}_{u,t}=O(\mathbf{h}_{u,t},\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k})$
where $\mathbf{h}_{u,t}$ is the hidden state at timestep $t$, $\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k}$ are the conditioning embeddings for different outcome types (e.g., $\mathbf{c}_{1}=\mathbf{e}_{\text{action}}$, $\mathbf{c}_{2}=\mathbf{e}_{\text{surface}}$), and $\hat{\mathbf{i}}_{u,t}$ is the resulting conditioned item representation.
Temporal Multi-token Prediction. Multi-token prediction in language models involves predicting multiple successive tokens at positions $t+1$ to $t+k$ given the current position $t$ (gloeckle2024better). This is typically implemented using separate output heads for each future position. However, in recommender systems, previous research (Pinnerformer22) suggests that the difference in time between actions is much more important than the position index alone, because a single position increment can correspond to very different time intervals for active users versus dormant users. To address this, we propose a time-based multi-token approach, in which tokens are generated for each time interval $\delta$ (e.g., generating recommendations for the user for the present, 30 seconds later, 1 minute later, and so on). Specifically, we condition our output head on the relative temporal distance from the current action, enabling a single parameterized head to generate predictions at various future time intervals:
(2) $\hat{\mathbf{i}}_{u,t+\delta}=O(\mathbf{h}_{u,t},\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k},\mathbf{e}_{\delta})$
where $\mathbf{e}_{\delta}$ is the learnable embedding for temporal offset $\delta\in\{1,2,\ldots,k\}$, and $\hat{\mathbf{i}}_{u,t+\delta}$ represents the predicted item at timestep $t+\delta$.
4.2. Input Sequence Representation
Given the heterogeneous nature of user interactions $(i_{j},a_{j},s_{j},t_{j})$ in our interaction history $H(u,t_{\text{max}})$, we need to transform each element into appropriate representations for the transformer architecture. A crucial design decision is defining representations for heterogeneous item types. There are two main ways to represent items in a sequence: (1) by using dense embedding vectors (HSTU24; KangM18), and (2) by learning a sequence of discrete IDs for each item, known as Semantic IDs (Tiger23; SemanticIDGeneral24). We explored both approaches and ultimately chose to use dense embedding vectors. In our experiments, we observed that Semantic IDs frequently suffered from representational collapse—many items were assigned the same Semantic ID, as also noted in (MultiBehaviorSASRec24; Hourglass24)—and consistently produced lower evaluation metrics (see Table 1(UnifyingGenRetrieval24). While optimizing performance with Semantic IDs remains an interesting direction for future work, our core contributions are representation-agnostic.
We now detail how each component of the interaction tuple $(i_{j},a_{j},s_{j},t_{j})$ is processed and integrated into our model architecture.
Item Embeddings. We generate dense embeddings for different entity types ($\tau\in\mathcal{T}$) using type-specific embedders that are learned end-to-end with the model. Each item type $\tau$ has a dedicated embedder $f_{\tau}$ that maps type-specific features to the shared representation space. While we describe embedders for Pins and search queries in this work, the approach can be extended to other entity types such as boards, users, or topics.
Pin Embeddings. For Pins, we combine semantic embedding features with learned Id embeddings (10.1145/3640457.3688053). The semantic representation uses the pre-trained OmniSage embedding (badrinath2025omnisage), which captures visual and textual content along with the engagement via the Pin-board graph. Additionally, we implement a large-scale ID embedding system where each Pin identifier is hashed into $k$ different hash functions, each mapping to embedding tables of size $d/k$. These $k$ embeddings are concatenated to form a $d$-dimensional ID embedding, providing a distributed representation that scales efficiently. The ID embeddings constitute sparse parameters that allow us to scale model capacity with minimal impact on serving latency and computational cost, as only a small subset of embeddings are accessed for any given batch of items. The concatenated OmniSage and ID embeddings are passed through 2-3 fully connected layers with ReLU activation and layer normalization, followed by $L_{2}$ normalization to produce the final Pin embedding $\mathbf{i}_{\text{pin}}$.
Query Embeddings. For search queries, we use only semantic representations as ID embeddings have not provided value in offline metrics for this entity type. We leverage the pre-trained OmniSearchSage query representation (agarwal2024omnisearchsage) which captures semantic query understanding. This representation is passed through a similar MLP architecture (2-3 fully connected layers with ReLU activation and layer normalization) followed by $L_{2}$ normalization to generate the final query embedding $\mathbf{i}_{\text{query}}$.
Temporal Embeddings. For each interaction timestep $t_{j}$, we extract absolute timestamp and time since previous action, capturing both global temporal context and local session dynamics. Our encoder applies sinusoidal transformations at multiple frequencies. For absolute timestamps, we use predefined periods $P=\{p_{1},p_{2},\ldots,p_{k}\}$ to capture daily, weekly, and seasonal patterns: $\sin(t_{\text{abs}}/p_{i}),\cos(t_{\text{abs}}/p_{i})$. For relative time differences, we use log-scale frequencies augmented with learnable phase parameters.
Sequence Construction. The final input sequence to the transformer is constructed by concatenating item and temporal embeddings, then adding surface embeddings for each interaction. This creates a rich contextual representation that captures what users interacted with (item), when they interacted (temporal), and where the interaction occurred (surface) within the platform.
4.3. Training Objective
During training, we employ teacher forcing with a next-item prediction objective. The hidden states $\mathbf{h}_{u,t}$ are processed through the conditioned output head $O$ to generate item representations that predict future items in the sequence. For outcome-conditioned generation, we condition the output head on the actual values of the target item during training. For multi-token prediction, we sample $k$ positions from the next $N$ positions in the sequence to serve as prediction targets.
Following Pinnerformer22, we compute similarity between the predicted item representation $\hat{\mathbf{i}}_{u,t}$ and candidate item $\mathbf{i}_{c}$ as $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{i}_{c})$, where $Q(\mathbf{i}_{c})$ is a bias correction term estimated using count-min sketches (cormode2005improved) to account for the non-uniform sampling of in-batch negatives.
The sampled softmax loss maximizes similarity to target items relative to negative samples $\mathbf{N}$:
(3) $L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}})=-\log\frac{\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}}))}{\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}}))+\sum_{\mathbf{i}_{n}\in\mathbf{N}}\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{n}))}$
We average this loss over all positions in each user sequence. For standard next-token prediction, the target is simply the immediate next item:
(4) $L_{\theta}^{\text{next}}(H(u,t_{\max}))=\frac{1}{|H(u,t_{\max})|}\sum_{t=1}^{t_{\max}-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
However, this strict sequential assumption may not align well with user interaction patterns. User interaction sequences differ fundamentally from language: while they exhibit global structure reflecting long-term interests, they lack the local structural constraints found in natural language. On social media platforms, users engage with feed suggestions in arbitrary order within sessions, making strict positional dependencies less meaningful than in natural language modeling.
To better accommodate these interaction patterns, we introduce an intra-feed relaxation that considers all items engaged within the same feed session as valid targets. Given a feed session $F_{s}=\{i_{t_{1}},i_{t_{2}},\ldots,i_{t_{n}}\}$ where $t_{1}<t_{2}<\ldots<t_{n}$, we define the relaxed target set for position $t_{j}$ as:
$T_{t_{j}}^{\text{feed}}=\{i_{t_{k}}\mid t_{k}\in F_{s},t_{k}>t_{j}\}$
This enables a relaxed objective that optimizes for predicting any future item within the same session:
(5) $L_{\theta}^{\text{feed}}(H(u,t_{\mathrm{max}}))=\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t_{j}}\min_{i\in T_{t_{j}}^{\text{feed}}}L_{s}(\hat{\mathbf{i}}_{u,t_{j}},i)$
where we minimize the loss to the best-matching item among all future items within the same feed session, rather than strictly requiring prediction of the immediate next item. Intuitively, this optimization can be posed as: can we predict any of the future items that the user will engage with within the same feed session? We find that this approach is more naturally aligned with user interaction patterns on Pinterest compared to strict next-token prediction.
Finally, we minimize the expected loss across all user sequences in dataset $\mathcal{D}$:
(6) $\arg\min_{\theta}\mathbb{E}_{H(u,t_{\max})\sim\mathcal{D}}[L_{\theta}^{\text{feed}}(H(u,t_{\max}))].$
4.4. Inference
Our PinRec model serves as a candidate generator within the broader recommendation stack, producing diverse item embeddings. The model operates in an autoregressive fashion to generate multiple candidate embeddings, which are then used to retrieve actual items through approximate nearest neighbor search.
To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer. The specific generation process varies based on the model configuration:
Unconditional Generation. For the baseline model, generation is straightforward: at each step, the model produces a single item embedding, which is appended to the sequence along with the target surface embedding and a predefined temporal offset (e.g., 10 seconds). This updated sequence is then used to generate the next embedding.
Outcome-Conditioned Generation. For models with outcome conditioning, for each outcome we generate a candidate embedding, producing multiple candidate embeddings per generation step. For autoregressive continuation, we sample a value for each condition according to the requested distribution over each condition $D_{1},D_{2},\ldots,D_{n}$ and use the corresponding embedding as input for the next step (Fig. 2).
Temporal Multi-Token Generation. When using temporal multi-token prediction, we sample time deltas up to a specified maximum and generate multiple embeddings at different temporal offsets, combined with any specified outcome conditions. For the next autoregressive input, we use the embedding corresponding to the immediate next token.
Across $n$ rollouts, this process generates varying numbers of embeddings depending on the model configuration and input settings. The generated embeddings are then used for candidate retrieval through Approximate Nearest Neighbor (ANN) search.
Budget Allocation. To retrieve $N$ total items, we first allocate retrieval budget across all generated embeddings. For unconditional models, the budget is distributed equally among all embeddings. For outcome-conditioned models, allocation is performed independently for each action type according to the specified budget constraints $\mathcal{B}$, ensuring that each condition receives its designated proportion of the total retrieval budget.
Embedding Compression. Given that autoregressive generation produces multiple output representations, we observe that many embeddings are often similar to one another. During ANN retrieval, these similar embeddings would collapse into overlapping sets of retrieved items, reducing diversity. To address this, we implement an embedding compression strategy: each generated embedding is merged into a previously uncompressed embedding (in generation order) if their cosine similarity exceeds a predefined threshold. When embeddings are compressed together, their allocated budgets are summed to determine the total retrieval budget for the compressed group.
This candidate generation process enables PinRec to serve as a flexible first-stage retrieval system within Pinterest’s recommendation infrastructure. The generated candidates can be further refined by downstream ranking models, allowing the system to benefit from both the generative model’s ability to capture complex user patterns and the precision of specialized ranking algorithms. The outcome-conditioning and multi-token capabilities make the system particularly suitable for multi-objective optimization scenarios common in production recommendation systems.